2x+5y=8
6x+y=10
Can you help me solve this question

Answer by rfer(16322)

y=-6x+10
2x+5y=8
2x+5(-6x+10)=8
2x-30x+50=8
-28x=-42
x=1.5
--------------
2*1.5+5y=8
3+5y=8
5y=5
y=1
-------------
6*1.5+1=10
9+1=10